# Modulating CNN Features with Pre-Trained ViT Representations for Open-Vocabulary Object Detection

### v1

@@ 3.5 VMC Module Design
Inspired by ViT-CoMer [30], a strong backbone network utilizing the pretrained ViT via feature interaction, we design the VMC module referring to its CNN-Transformer bidirectional fusion interaction(CTI) block. As shown in Figure 2, VMC module mainly consists of a modulating information generation (MIG) block and feature modulation (FM) blocks. Notice that there are several crucial differences between ViT-CoMer and our design.
First, VMC module focuses on injecting the information from a sophisticated visual encoder into CNN features, which is unidirectional instead of bidirectional. To keep the rich knowledge in visual encoder, the vision transformer in VMC is frozen while ViT-CoMer is fine-tunable. Moreover, VMC module is applied merely at the end of two branches, which owns different structure from it. Our method is designed for OVOD task, while ViT-CoMer aims to enhance the closed-set object detector.
Modulating Information Generation. We attempt to utilize the ViT features collected from different blocks to generate a unified representation. Thus, we propose a novel module named as modulating information generation. Its structure is illustrated in Figure 3. Firstly, the ViT features , and are concatenated along the channel dimension. This concatenated high-dimensional features contains representation from different depth in ViT. Then, a linear projection layer is applied, which not only reduce the channel number from to , but also simply adapting the ViT features to the downstream processing. We implement this linear projection with a fully connected layer. Via MIG, we could attain the adaptive modulating feature. The processing of MIG could be denoted as:
Feature Modulation. With the adaptive modulating features from MIG and the multi-scale features from CNN branch, we achieve the feature modulation by stacking a serial of feature modulation (FM) blocks. As shown in Figure 2, FM block takes the multi-scale features and the modulating feature as inputs. The output of the last block is sent to the next in a manner of cascaded refinement. All blocks use the same modulating feature .
FM shares the same structure as ViT-to-CNN (VtoC) block in ViT-CoMer, its structure is illustrated in Figure 4. This block consists of a Multi-scale deformable attention (MSDA), two layer normalization operations, two feed-forward networks and three skip connection sums.
@@
Whereas the element-wise sum operation simply fuses the features, the representation from ViT is merely injected into the features at second scale. As we know, deformable attention (DA) utilizes the reference points to sample the features to attain the values. In MSDA, each scale will provide a fixed number of reference points, that is, ViT representation could be scattered into the features at all scales via MSDA. The output of MSDA is then processed similarly as the above:
Through the operations in first FM, features at each scale of has carried the representations from frozen ViT. In the next FM block, we repeat the processing as illustrated above to further inject the ViT representations into the multi-scale features. For -th block, feature modulation is denoted as:
We also study the choice of in Section 4.3.1. The output of final block is denoted as , we interpolate the feature with largest scale in and add it with . After batch normalization, this four-scale features could be used as the backbone features for downstream processing.
@@ 4.1.1 Datasets and evaluation metrics
OV-COCO. MSCOCO is one of the most commonly used datasets in object detection. To evaluate a detector under open-vocabulary setting, OVR-CNN [32] proposes to build the base set with 48 categories of its data, and the other 17 categories of data are used as the novel. We follow this rule to train our method with the base set data and evaluate it with the set including both the base and the novel. As the behaviors of existing works, bounding box AP at IoU threshold 0.5 of novel categories is employed as the main metric, which denoted as AP.
OV-LVIS. LVIS is a large-scale dataset for instance segmentation, which contains 1203 categories of objects. The categories is split into rare (337 categories), common (461) and frequent(405) based on the number of images the they appear. We follow ViLD [4] to split the dataset, rare categories are split into the novel while common and frequent categories are split into the base. The mask AP averaged on IoUs from 0.5 to 0.95 of rare categories, denoted as mAPr, is chosen as the main metric for evaluation.
@@ 4.1.2 Implementation Details
Our detection framework is based on F-ViT [27]. We use the self-distilled visual encoder from CLIPSelf to initialize the weights of vision transformer branch. For OV-COCO, we use the CLIPSelf with proposal distillation setting while OV-LVIS is the patch distillation setting. We conduct the experiments with 8 RTX 3090 GPUs and set the batch size as 1 on each GPU. For fair comparisons, our method inherits the configurations of baseline except the proposed part. Following the implementation of F-ViT, we use the same training scheme. Utilizing AdamW optimizer with a learning rate of , detector for OV-COCO is trained for 3 epochs while the one for OV-LVIS is 48-epoch training.
@@ 4.2 Quantitative Comparisons
F-ViT+VMCNet | VMCNet-L | 48.5 (+4.0) |
OV-COCO. We report the comparison with previous methods in Table 1. VMCNet based on ViT-B/16 is denoted as VMCNet-B while VMCNet-L refers to the one based on ViT-L/14. To get rid of the influence from experiment environment, we run the baseline method under our local environment and display the results. Under ViT-B/16 setting, our method achieves 44.3 AP, which is even comparable to the baseline with ViT-L/14. The performance can be further improved when increasing the scale of network. F-ViT equipped with VMCNet-L achieves 48.5 AP and also outperforms the baseline method by an obvious margin of 4.0%. Through the evaluation on OV-COCO, we found VMCNet could effectively boost the detection performances on novel categories.
OV-LVIS. Table 2 lists the state-of-the-art methods on OV-LVIS. VMCNet also surpasses previous state of the arts. Compared to the baseline with ViT-B/16, our method can bring +2.0 mAPr gain. Under ViT-L/14 setting, VMCNet can still outperform the baseline approach by 1.5 mAPr. Therefore,VMCNet is proved to be effective on detecting objects of novel categories.
@@
F-ViT+CLIPSelf | ViT-B/16 | 25.8 |
F-ViT+VMCNet | VMCNet-B | 27.8(+2.0) |
F-ViT+CLIPSelf | ViT-L/14 | 36.9 |
F-ViT+VMCNet | VMCNet-L | 38.4(+1.5) |
@@ 4.3.1 Effectiveness of model structure
The ablation study on the main components is shown in Table 3. We start from the baseline method which only applies the CLIP ViT, its performance is reported in first row. In second row, the last few transpose convolutional layers to interpolate the ViT features are removed and backbone is replaced by the trainable CNN, where ViT only serves as the RoI classifier. This design corresponds to the illustration in Figure 1(b), where detector could only rely on the features extracted by the simple CNN and thus have poor performance. In third row, we further add the abridged FM blocks which does not add , though the performance is improved by a little, detector still suffers from lack of ViT representation. In the bottom line, the entire VMCNet is implemented. We can see that the increase on parameter cost is lower than 4M. Therefore, these main components can effectively utilize the information from both the frozen ViT and base training data.
@@
Besides, we explore the strategy of placing FM in VMC modules. As shown in Table 4, we presents 5 strategies. We can see that the highest AP is achieved in the second row, that is, the optimal value of is 3. To better understand these strategies, the local designs in 2-nd, 4-th and 5-th row are illustrated in Figure 5(a), (b) and (c), respectively. Compared to the 2-nd row, the 4-th adds a repeated structure behind, which does not improve performance. In 5-th line, each FM block is supported by a different MIG module, its result shows using more MIG modules can not bring gain, either. Therefore, we choose the strategy in second row for OV-COCO setting.
We also analyze the impact of the number of MRFP modules in CNN branch. The first row in Table 5 corresponds to the structure removing MRFP from VMCNet, which leads a drop of 1.4 AP. The bottom line is the result of cascading two MRFP modules, which shows that increasing does not improve the performance. Thus, only one MRFP module is used in our CNN branch design.
@@ 4.3.2 Effectiveness of ViT branch
In this part, we conduct ablation experiments on the used ViT feature layers as shown in Table 6. The default setting we use is the 3-rd column. In first three rows, one of the default feature layers is removed. As a result, these results are inferior to the default setting in the 4-th line. However, in fifth row, the detection performance on novel categories degenerates obviously when the extra ViT layer is added. In the last row, features from the deep blocks of ViT are applied, which leads to the poorer performance. Thus, we find that the ability to detect novel categories may benefit from features of some certain layers of frozen ViT. In OV-COCO setting, features from the high level of ViT may even hurt the performances.
@@ 4.4 Qualitative Visualization
In Figure 6, we visualize the detection results of the baseline method and ours. These image samples are collected from COCO validation set. From the visual comparisons, we observe that our method is more confident to recognize the novel objects. For example, in second column, the baseline fails to detect the bus in image, while ours could predict it with a relatively high confidence score.
@@ 5 Conclusion
In this paper, we propose a two-branch backbone network for OVOD task, named as VMCNet. Our method combines the advantages of CNN and ViT. The frozen ViT branch saves the generalization ability while the trainable CNN learns information of the base training data. At last, the proposed VMC module could modulate multi-scale convolutional features with the representations from ViT branch. Our design provides an effective scheme to utilize the knowledge from both the pre-training and base training. As a result, VMCNet improves the detection performances on novel categories effectively and efficiently. Experimental results demonstrate that our backbone network outperforms the baseline and reaches the level of state-of-the-art methods.

### v2

@@ 3.5 VMC Module Design
Inspired by ViT-CoMer [32], a strong backbone network utilizing the pretrained ViT via feature interaction, we design the VMC module based on its CNN-Transformer bidirectional fusion interaction (CTI) block. As shown in fig. 2, VMC module mainly consists of a modulating information generation (MIG) block and feature modulation (FM) blocks. Notice that there are several crucial differences between ViT-CoMer and our method.
First, VMC module focuses on injecting the information from a sophisticated visual encoder into CNN features, which is unidirectional instead of bidirectional. To keep the rich knowledge in visual encoder, the vision transformer in VMCNet is frozen while ViT-CoMer is fine-tunable. Moreover, VMC module is applied merely at the end of two branches, which owns different structure from theirs. Our method is designed for OVOD task, while ViT-CoMer aims to enhance the closed-set object detector.
Modulating Information Generation. We attempt to utilize the ViT features collected from different blocks to generate a unified representation. Thus, we propose a novel module named as modulating information generation. Its structure is illustrated in fig. 3. Firstly, the ViT features , and are concatenated along the channel dimension. The concatenated high-dimensional features contains representations from different depths in ViT. Then, a linear projection layer is applied, which not only reduces the channel number from to , but also simply adapts the ViT features to the downstream processing. We implement this linear projection with a fully connected layer. Via MIG, we could attain the adaptive modulating feature. The processing of MIG could be denoted as:
Feature Modulation. With the adaptive modulating features from MIG and the multi-scale features from CNN branch, we achieve the feature modulation by stacking a serial of feature modulation (FM) blocks in a cascaded manner as shown in fig. 2. Output of the previous FM block is sent to the next one to achieve progressive feature refinement. All blocks use the same modulating feature .
FM block shares the same structure as ViT-to-CNN (VtoC) block in ViT-CoMer, its structure is illustrated in fig. 4. This block consists of a multi-scale deformable attention (MSDA) [42], two layer normalization (LN) operations, two feed-forward networks (FFN), two dropout layers and three skip connection sums.
@@
Although the element-wise sum operation simply fuses the features, the representation from ViT is merely injected into the features at second scale. As we know, deformable attention (DA) utilizes the reference points to sample the features to attain the values. In MSDA, each scale will provide a fixed number of reference points, that is, ViT representation could be scattered into the features at all scales via MSDA. The output of MSDA is then processed similarly as the above:
Through the operations in first FM block, features at each scale of has carried the representations from frozen ViT. In the next block, we repeat the processing as illustrated above to further inject the ViT representations into the multi-scale features. For -th () block, feature modulation is denoted as:
We also study the choice of in section 4.3.1. The output of final block is denoted as , we interpolate the feature with largest scale in and add it with . After batch normalization, the four-scale features could be used as the backbone features for downstream steps. The processing for is represented as below:
@@ 4.1.1 Datasets and evaluation metrics
OV-COCO. MSCOCO is one of the most commonly used datasets in object detection. To evaluate a detector under open-vocabulary setting, OVR-CNN [34] proposes to build the base set with 48 categories of its data, and the other 17 categories of data are used as the novel. We follow this rule to train our method with the base set data and evaluate it with the set including both the base and the novel set. As the behaviors of existing works, bounding box AP at IoU threshold 0.5 of novel categories is employed as the main metric, which is denoted as AP.
OV-LVIS. LVIS is a large-scale dataset for instance segmentation, which contains 1203 categories of objects. The category set is split into rare (337 categories), common (461) and frequent (405) based on the number of images the objects appear. We follow ViLD [4] to split the dataset. Rare categories are split into the novel while common and frequent categories are split into the base. The mask AP averaged on IoUs from 0.5 to 0.95 of rare categories, denoted as mAPr, is chosen as the main metric for evaluation.
@@ 4.1.2 Implementation Details
Our detection framework is based on F-ViT [29]. We use the self-distilled visual encoder from CLIPSelf to initialize the weights of vision transformer branch. For OV-COCO, we use the CLIPSelf with proposal distillation setting and leverage , and . For OV-LVIS, CLIPSelf with the patch distillation setting is used and , , , and are employed. We conduct the experiments with 8 RTX 3090 GPUs and set the batch size as 1 on each GPU. For fair comparisons, our method inherits the configurations of baseline except the proposed part. Following the implementation of F-ViT, we use the same training scheme. AdamW optimizer with a learning rate of is utilized, detector for OV-COCO is trained for 3 epochs while 48-epoch training is for OV-LVIS.
@@ 4.2 Quantitative Comparisons
F-ViT+VMCNet (Ours) | VMCNet-B | 44.3 (+5.6) |
F-ViT+VMCNet (Ours) | VMCNet-L | 48.5 (+4.0) |
OV-COCO. We report the comparison with previous methods in table 1. VMCNet based on ViT-B/16 is denoted as VMCNet-B while VMCNet-L refers to the one based on ViT-L/14. To get rid of the influence from experiment environment, we train the baseline method under our local environment and display the results. We choose the highest performance of baseline method for comparison. Under ViT-B/16 setting, our method achieves 44.3 AP, which is even comparable to the baseline with ViT-L/14. The performance can be further improved when increasing the scale of network. F-ViT equipped with VMCNet-L achieves 48.5 AP and also outperforms the baseline method by an obvious margin. Through the evaluation on OV-COCO, we found VMCNet could effectively boost the detection performances on novel categories.
OV-LVIS. Table 2 lists the performances of state-of-the-art methods on OV-LVIS. VMCNet also surpasses the previous methods. Compared to the baseline with ViT-B/16, our method can bring +2.3 mAPr gain. Under ViT-L/14 setting, VMCNet can outperform the baseline approach by 3.2 mAPr. Therefore, VMCNet is proved to be effective on detecting objects of novel categories.
@@
F-ViT+CLIPSelf [29] | ViT-B/16 | 25.5 |
F-ViT+CLIPSelf [29] | ViT-L/14 | 35.2 |
F-ViT+VMCNet (Ours) | VMCNet-B | 27.8(+2.3) |
F-ViT+VMCNet (Ours) | VMCNet-L | 38.4(+3.2) |
@@ 4.3.1 Effectiveness of model structure
The ablation study on the main components is shown in table 3. We start from the baseline method which only applies the CLIP ViT, its performance is reported in first row. In second row, backbone is replaced by the trainable CNN, ViT only serves as the RoI classifier and the last few transposed convolutional layers to interpolate its features are removed. This design corresponds to the illustration in fig. 1(b), where detector could only rely on the features extracted by the simple CNN and thus have poor performance. In third row, we further add the abridged FM blocks which remove the addition, though the performance is improved by a little, detector still suffers from lack of ViT representation. In the bottom line, the entire VMCNet is implemented. We can see that the increase on parameter cost is lower than 4M. Therefore, these main components can effectively utilize the information from both the frozen ViT and base training data.
@@
Besides, we explore the strategy of placing FM blocks in VMC modules. As shown in table 4, we present 5 strategies. We can see that the highest AP is achieved in the second row, that is, the optimal value of is 3. To better understand these strategies, the local designs in 2-nd, 4-th and 5-th row are illustrated in fig. 5(a), fig. 5(b) and fig. 5(c), respectively. Compared to the 2-nd row, the 4-th adds a repeated structure behind, which does not improve performance. In 5-th line, each FM block is supported by a different MIG module. This result shows using more MIG modules can not bring gain, either. Therefore, we choose the strategy in second row for OV-COCO setting.
We also analyze the impact of the number of MRFP modules in CNN branch. The first row in table 5 corresponds to the structure removing MRFP from VMCNet, which leads to a drop of 1.4 AP. The bottom line is the result of cascading two MRFP modules, which shows that increasing does not improve the performance. Thus, only one MRFP module is used in our CNN branch design.
@@ 4.3.2 Effectiveness of ViT branch
In this part, we conduct ablation experiments on the used ViT feature layers as shown in table 6. The default setting we use is the 4-th row. In first three rows, one of the default feature layers is removed. As a result, these results are inferior to the setting in 4-th line. However, in 5-th row, the detection performance on novel categories degenerates obviously when the extra ViT feature layer is added. In the last row, features from the deep blocks of ViT are applied, which leads to the poorer performance. Thus, we find that the ability to detect novel categories may benefit from features of some certain layers of frozen ViT. In OV-COCO setting, features from the high level of ViT may even hurt the performances.
@@ 4.4 Qualitative Visualization
Detection results of the baseline method and ours are visualized in fig. 6. We set the score threshold as 0.3 to filter the predicted bounding box whose classification score is lower than this value. From the visual comparisons, we observe that our method is more confident to recognize the novel objects. For example, in second column, the baseline fails to detect the bus in image, while ours could predict it with a relatively high confidence score.
@@ 5 Conclusion
In this paper, we propose a two-branch backbone network for OVOD task, named as VMCNet. Our method combines the advantages of CNN and ViT. The frozen ViT branch saves the generalization ability while the trainable CNN learns information of the base training data. At last, the proposed VMC module could modulate multi-scale convolutional features with the representations from ViT branch. Our design provides an effective scheme to utilize the knowledge from both the pre-training and base training. As a result, VMCNet improves the detection performances on novel categories effectively and efficiently. Experimental results demonstrate that our method improves the baseline significantly and outperforms the state-of-the-art methods.